# Disentangling the Complex Multiplexed DIA Spectra in De Novo Peptide Sequencing

### v5

@@ Appendix A Additional Features for Spectrum
    Local Rank: Similar to the global rank, but restricted to peaks within the ±$\pm 50$ Th m/z range, providing a localized perspective on peak significance. Computed as $\frac{R_{local}}{N_{local}}$, similar to rank.
@@ Appendix B Ion Offsets
List of supportingtables 1 Ion offsets for ion types we used. [N] is the molecular mass of the neutral N-terminal group; [C] is the molecular mass of the neutral C-terminal group. C,H,O are the mass of the carbon atom, hydrogen atom and oxygen atom individually. [13]
Appendix C Type of Labeled Ions in Pretraining
@@
Appendix D Precursor Feature Detection
For the detection of precursor features from liquid chromatography-mass spectrometry (LC-MS) maps, we utilized the same standard set of precursor information as employed by DeepNovo-DIA [19]. In our experiments, we implemented the detection results from DIA-NN [29] during the main experiment, and DIAUmpire [45] during the comparison with Cascadia; however, these can be substituted with outputs from other existing peak detection algorithms, such as those referenced in Zhang et al. (2012) [38], Taynova et al. (2016) [46] or Tsou et al. (2015) [45]. The outcome of this detection step is a list of precursor features, each comprising the following essential information: feature ID, precursor mass-to-charge ratio (m/z), charge state, retention-time center, and scans across the retention-time range.
@@ Appendix D Precursor Feature Detection
Appendix E Characteristics of Test Datasets
@@
Dataset Coeluting Number # Signal Peaks # Noise Peaks Median Peptide Length Median Noise Intensity Isolation Window Mass Spectrometer
@@
Appendix F Links to Datasets
Supporting Table 4 provides links to our training and testing datasets. Every dataset is searched with DIA-NN 1.8.1 [29] with carboxymethylation of cysteine as fixed modification, and oxidation of methionine as variable. PSMs under 1% FDR is kept for training or testing. The raw files are converted to mzML with msConvert 3.0 with vendor peak picking. Then the PSMs and mzMLs are converted to our customized format with our processing codes, by extracting the related spectrums of each PSM. For de novo only analysis, feature detection is done on the mzML files with DIA-Umpire 2.2.8 [45].
@@ Appendix F Links to Datasets
Appendix G Model Implementation Details
Our model is trained on a Lion optimizer [49] with learning rate $2\times 10^{-6}$, over 3 epochs, with early stop based on validation error. The model includes 4 layers both on encoder and decoder side, with 1,024 hidden size and 8 attention heads.
For older-generation data, training and validation sets are Pain, PXD019777, and PXD003179, with 680,947/254,467/1,206,052 PSMs respectively. For Astral data, they are PXD046386 or PXD046453 + PXD046453, with 1,896,662, 1,086,577, and 2,136,215 PSMs respectively, these PSMs are chosen from the DIA-NN search result with 1% FDR.
To prevent information leakage and ensure a rigorous evaluation of model generalization, we designed our training/validation/test splits to fully separate related peptide signals across sets, including coeluting peptides. Our model architecture leverages not only the target peptide but also contextual information from coeluting peptides within a spectrum. Therefore, to eliminate any direct or indirect sequence overlap, we adopted a raw file–based splitting strategy: for each training run, we designated a specific RT range in one raw file exclusively for validation and used the remaining for training. This guarantees that no peptides, or their coeluting partners, appear in both training and validation sets.
@@ Appendix G Model Implementation Details
For the test set, we ensured strict non-overlap in both peptide sequences and experimental conditions. All test sequences were drawn from datasets entirely distinct from those used in training. Furthermore, in our cross-species evaluation using Orbitrap Astral data, we imposed an even stricter criterion: training and testing data come from different species. Specifically, when training on the human datasets PXD046453 + PXD046444, we tested only on the yeast dataset PXD046386. Conversely, when training on the yeast dataset PXD046386, we used entirely separate human datasets for testing.
Appendix H Visualization of Learned Embeddings
@@
We generate a t-SNE [50] plot from the learned peak embeddings, and observe that the model implicitly captures the relationship between coeluting precursors and their corresponding fragment ions, even when trained solely with the ion type loss. The t-SNE plots (shown in Supporting Figure 1, 2, and 3) further reveal that peak embeddings are organized not only by fragment ion type (ion type label) but also by their source peptide (ion source label). For instance, in Supporting Figure 1, a dark blue cluster on the right side of the graph represents peaks originating from peptide 1 in the ion source plot. Within this cluster, the peaks are further subdivided into distinct colors (pink and green) in the ion type plot, corresponding to different fragment ion types. We provide the t-SNE plots from three different peptides to illustrate this effect.

### v6

@@ Appendix C Additional Features for Spectrum
    Local Rank: Similar to the global rank, but restricted to peaks within the $\pm 50$ Th m/z range, providing a localized perspective on peak significance. Computed as $\frac{R_{local}}{N_{local}}$, similar to rank.
@@ Appendix D Ion Offsets
List of supportingtables 1 Ion offsets for ion types we used. [N] is the molecular mass of the neutral N-terminal group; [C] is the molecular mass of the neutral C-terminal group. C,H,O are the mass of the carbon atom, hydrogen atom and oxygen atom individually. (14)
Appendix E Type of Labeled Ions in Pretraining
@@
Appendix F Precursor Feature Detection
For the detection of precursor features from liquid chromatography-mass spectrometry (LC-MS) maps, we utilized the same standard set of precursor information as employed by DeepNovo-DIA (20). In our experiments, we implemented the detection results from DIA-NN (27) during the main experiment, and DIAUmpire (43) during the comparison with Cascadia; however, these can be substituted with outputs from other existing peak detection algorithms, such as those referenced in Zhang et al. (2012) (36), Taynova et al. (2016) (44) or Tsou et al. (2015) (43). The outcome of this detection step is a list of precursor features, each comprising the following essential information: feature ID, precursor mass-to-charge ratio (m/z), charge state, retention-time center, and scans across the retention-time range.
@@ Appendix F Precursor Feature Detection
Appendix G Characteristics of Test Datasets
@@
Dataset
Coeluting
Number
# Signal
Peaks
# Noise
Peaks
Median
Peptide
Length
Median
Noise
Intensity
Isolation
Window
Mass
Spectrometer
@@
Appendix H Links to Datasets
Supporting Table 4 provides links to our training and testing datasets. Every dataset is searched with DIA-NN 1.8.1 (27) with carboxymethylation of cysteine as fixed modification, and oxidation of methionine as variable. PSMs under 1% FDR is kept for training or testing. The raw files are converted to mzML with msConvert 3.0 with vendor peak picking. Then the PSMs and mzMLs are converted to our customized format with our processing codes, by extracting the related spectrums of each PSM. For de novo only analysis, feature detection is done on the mzML files with DIA-Umpire 2.2.8 (43).
@@ Appendix H Links to Datasets
Appendix I Model Implementation Details
Our model is trained on a Lion optimizer (47) with learning rate $2\times 10^{-6}$, over 3 epochs, with early stop based on validation error. The model includes 4 layers both on encoder and decoder side, with 1,024 hidden size and 8 attention heads.
For older-generation data, training and validation sets are Pain, PXD019777, and PXD003179, with 680,947/254,467/1,206,052 PSMs respectively. For Astral data, they are PXD046386, PXD046453, and PXD046444, with 1,896,662, 1,086,577, and 2,136,215 PSMs respectively, these PSMs are chosen from the DIA-NN search result with 1% FDR.
@@ Appendix I Model Implementation Details
For the test set, we ensured strict non-overlap in both peptide sequences and experimental conditions. All test sequences were drawn from datasets entirely distinct from those used in training. For example, when training on the human datasets PXD046453 + PXD046444, we tested only on the yeast dataset PXD046386. Conversely, when training on the yeast dataset PXD046386, we used entirely separate human datasets for testing.
Appendix J Sensitivity to Coelution Number
Appendix K Ablation Study
In this ablation study, illustrated in Figure 2, we evaluate the impact of different configurations on our model using the Yeast KO dataset. The removal of coelution-aware pretraining reduces peptide recall by 8.5% compared to DIANovo. The removal of Rotary Positional Encoding (RoPE) results in peptide recall decreasing to 67.2% of the original. Excluding extended peak features or limiting the input to only three neighboring spectra leads to moderate decreases in performance. Finally, binning spectra into 0.5 Th intervals produces a substantial reduction in performance.
((a)) Relative amino acid recall
((b)) Relative peptide recall
((c)) Relative amino acid precision
List of supportingfigures 2 Relative amino acid recall (a), peptide recall (b), and amino acid precision (c) of our method vs different configurations, compared to our method, on the Yeast KO dataset.
Appendix L Visualization of Learned Embeddings
@@
We generate a t-SNE (48) plot from the learned peak embeddings, and observe that the model implicitly captures the relationship between coeluting precursors and their corresponding fragment ions, even when trained solely with the ion type loss. The t-SNE plots (shown in Supporting Figure 3, 4, and 5) further reveal that peak embeddings are organized not only by fragment ion type (ion type label) but also by their source peptide (ion source label). For instance, in Supporting Figure 3, a dark blue cluster on the right side of the graph represents peaks originating from peptide 1 in the ion source plot. Within this cluster, the peaks are further subdivided into distinct colors (pink and green) in the ion type plot, corresponding to different fragment ion types. We provide the t-SNE plots from three different peptides to illustrate this effect.